OpenStudy (anonymous):

How do you convert 0.1232323... into a fraction? Please explain

OpenStudy (anonymous):

Let x=0.1232323.... 10 x=1.232323... (1) 100 x=12.32323... 1000 x=123.2323... (2) (2)-(1) gives 990 x=122 x=122/990=?

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw